A steel ball rolls with a constant velocity across a table top 0.808 m high. It rolls off and hits the ground +0.264 m horizontally from the edge of the table. How fast was the ball rolling?
1
Expert's answer
2020-10-12T07:27:29-0400

Let's first write the equations of motion of the steel ball in horizontal and vertical directions:


"x = vt,""y = \dfrac{1}{2}gt^2,"

here, "x = 0.264 \ m" is the horizontal displacement of the ball, "v" is the ball's velocity across the table, "t" is time that the ball takes to reach the ground, "y = 0.808 \ m" is the vertical displacement of the ball (or the height of the table) and "g = 9.8 \ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's find the time that the ball takes to reach the ground from the second equation:


"t = \sqrt{\dfrac{2y}{g}}."

Then, we can substitute "t" into the first equation and find "v":


"x = v\sqrt{\dfrac{2y}{g}},""v = \dfrac{x}{\sqrt{\dfrac{2y}{g}}} = \dfrac{0.264 \ m}{\sqrt{\dfrac{2 \cdot 0.808 \ m}{9.8 \ \dfrac{m}{s^2}}}} = 0.65 \ \dfrac{m}{s}."

Answer:

"v = 0.65 \ \dfrac{m}{s}."